TWO SIGHTLINES TOWARD THE GALACTIC CENTER WITH REMARKABLE H_{3}^{+} and CO SPECTRA

\underline{TAKESHI OKA}, Department of Astronomy and Astrophysics, and Department of Chemistry, the University of Chicago, IL 60637; T. R. GEBALLE, Gemini Observatory, Hilo, HI 96720.

Until last year the known stars in the Galactic center with sufficiently smooth spectra and of sufficient brightness to be suitable for high resolution infrared absorption spectroscopy of H_{3}^{+} and CO were limited to a narrow range of longitudes, from the central cluster of hot stars to approximately 30 pc east of the center. In that region we have discovered a vast amount of warm (T ∼ 250 K) and diffuse (n ∼ 100 cm^{−3}) gas.^{a}^{b}. In order to more fully characterize the gas within the r∼ 180 pc Central Molecular Zone (CMZ) it is necessary to find additional such stars that cover a much wider longitudinal range of sightlines. We are in the process of identifying luminous dust-embedded objects suitable for spectroscopy between 170 pc east and west and 14 pc north and south of Sgr A* using the Spitzer GLIMPSE and the 2MASS catalogues.

Here we present spectra of H_{3}^{+} and CO towards two such objects, one located 140 pc west of Sgr A* close to Sgr E, and the other located on a line of sight to the Sgr B molecular cloud complex 85 pc to the east of Sgr A*.^{c} The sightline to the west passes through two dense clouds of unusually high negative velocities in Sgr E which show strong absorptions by both H_{3}^{+} and CO. This sightline also samples a portion of the Expanding Molecular Ring (EMR) which shows strong H_{3}^{+} absorption but undetectable CO clearly indicating that the gas is diffuse, similar to sightlines closer to Sgr A*. The H_{3}^{+} spectrum toward Sgr B reveals at least ten absorption components covering over 200 km s^{−1} and by far the largest equivalent width ever observed in an interstellar H_{3}^{+} line. The CO spectrum is similarly complex but covers a somewhat narrower velocity range indicating mixture of dense and diffuse environments is present along this line of sight. Together the CO and H_{3}^{+} spectra appear to provide the first near-infrared view into that hotbed of star formation.

Although further characterization of gas in those two sightlines and many more new sightlines must await observations of H_{3}^{+} and CO this July at the Gemini South Observatory, clearly we are opening up a new vista of the CMZ.

aOka, T, Geballe, T. R., Goto, M., Usuda, T., & McCall, B. J. 2005, ApJ, 632, 882

bGoto, M, et al. 2008, ApJ, 668, 306

cGeballe, T. R. & Oka, T. 2010, ApJL, 709, L70